OpenStudy (anonymous):

WILL MEDAL AND FAN The equation of the line of best fit of a scatter plot is y = –7x − 2. What is the the y-intercept? –7 –2 2 7

OpenStudy (epoweritheta):

in y=mx+c , c is the y intercept

OpenStudy (anonymous):

are you sure?

OpenStudy (epoweritheta):

yeah

OpenStudy (anonymous):

the equation is y=-7x-2, no addition is in the problem

OpenStudy (epoweritheta):

so y intercept is -2

OpenStudy (epoweritheta):

y= -7x+(-2)

OpenStudy (anonymous):

so now its B?

OpenStudy (anonymous):

OHHHHH NVM

OpenStudy (anonymous):

I GOT IT

OpenStudy (anonymous):

Can you possibly help with another one?

OpenStudy (epoweritheta):

yes

OpenStudy (anonymous):

What is the equation of this line of best fit in slope-intercept form? y = negative one halfx + 8 y = –2x + 8 y = –8x + one half y = –8x + 2

OpenStudy (anonymous):

sorry I will give you the answers again

OpenStudy (anonymous):

y=-1/2x+8 y=-2x+8 y=-8x+1/2 y=-8x+2

OpenStudy (anonymous):

@picgal

OpenStudy (anonymous):

@theATM

OpenStudy (anonymous):

@danny131

OpenStudy (anonymous):

@FluffyPancakeTerd

OpenStudy (anonymous):

@maddiegirl

OpenStudy (anonymous):

@MSxSammiBooxX

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@TwilightRain1

OpenStudy (anonymous):

@TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1 @TwilightRain1

OpenStudy (anonymous):

@LittleRedRiding

OpenStudy (anonymous):

y has i been tagged a gazillion times?